A train is traveling at 200mph, leaves station A at 1.PM. Train B, traveling at 300mph, leaves station B at 3.PM.Both stations are directly across from each other and X miles away. If the trains meet at 4.PM, what is X
A)900 mi
B)700 mi
C) 600 mi
D)400 mo

Answer by ikleyn(53765)

.

Before to meet each other, train A traveled 200 + 200 + 200 = 3*200 = 600 miles in 3 hours, 
while train B traveled only 300 miles in 1 hour.


These distances are in opposite directions from the meeting point.


Therefore, you need to add 600 miles and 300 miles to get your ANSWER.